# On homomorphic images of ultraproducts

### v1

@@ 4. A common element below two ultrafilters
Suppose that $\mathcal{U}\leq_{RK}\mathcal{U}^{\prime}$. Then every homomorphic image of an ultraproduct using $\mathcal{U}^{\prime}$ is also a homomorphic image of an ultraproduct using $\mathcal{U}$.
@@
The assertion Near coherence of filters (NCF) states that for any two nonprincipal ultrafilters $\mathcal{U}$ and $\mathcal{U}^{\prime}$ there exists a finite-to-one function $f:\omega\rightarrow\omega$ such that $f(\mathcal{U})=f(\mathcal{U}^{\prime})$ [2, §5]. Particularly, for any two nonprincipal $\mathcal{U}$ and $\mathcal{U}^{\prime}$, NCF gives an ultrafilter $\mathcal{U}^{\prime\prime}$ which is $\leq_{RK}$ below both $\mathcal{U}$ and $\mathcal{U}^{\prime}$. By Proposition 25 the homomorphic images of ultraproducts over $\mathcal{U}$ or $\mathcal{U}^{\prime}$ are also homomorphic images of ultraproducts over $\mathcal{U}^{\prime\prime}$. Since NCF is consistent with ZFC [3], Question 5 has a consistent positive answer.

### v2

@@ 4. A common element below two ultrafilters
Suppose that $\mathcal{U}\leq_{RK}\mathcal{U}^{\prime}$. Then every homomorphic image of an ultraproduct over $\mathcal{U}^{\prime}$ is also a homomorphic image of an ultraproduct over $\mathcal{U}$.
@@
The assertion Near coherence of filters (NCF) states that for any two nonprincipal ultrafilters $\mathcal{U}$ and $\mathcal{U}^{\prime}$ there exists a finite-to-one function $f:\omega\rightarrow\omega$ such that $f(\mathcal{U})=f(\mathcal{U}^{\prime})$ [2, §5]. In particular, for any two nonprincipal ultrafilters $\mathcal{U}$ and $\mathcal{U}^{\prime}$, NCF gives a nonprincipal ultrafilter $\mathcal{U}^{\prime\prime}$ which is $\leq_{RK}$ below both $\mathcal{U}$ and $\mathcal{U}^{\prime}$. By Proposition 26 the homomorphic images of ultraproducts over $\mathcal{U}$ or $\mathcal{U}^{\prime}$ are also homomorphic images of ultraproducts over $\mathcal{U}^{\prime\prime}$. Since NCF is consistent with ZFC [3], Question 5 has a consistent positive answer.
Acknowledgement
The author thanks the referees for pointing out clarifications and typos in a very timely manner.